OpenStudy (anonymous):

HELP PLEASE !!! 4X+5Y=13 3X+Y=-4

OpenStudy (anonymous):

@agent0smith

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@MERTICH

OpenStudy (anonymous):

please show steps

OpenStudy (anonymous):

What's the question

OpenStudy (johnweldon1993):

4x + 5y = 13 3x + y = -4 System of equations... Lets solve the bottom equation for 'y' 3x + y = -4 Subtract 3x from both sides to isolate 'y' 3x + y =-4 -3x -3x ------------ y = -3x - 4 Now take that...and substitute it into the first equation for every 'y' we see 4x + 5y = 13 would then become 4x + 5(-3x - 4) = 13 Distribute the 5... 4x - 15x - 20 = 13 Combine like terms... -11x - 20 = 13 Add 20 to both sides of the equation -11x = 33 divide both sides by -11... x = -3 Now...since we know x = -3....we can solve for 'y' by plugging that into either of our original equations 4x + 5y = 13 becomes 4(-3) + 5y = 13 -12 + 5y = 13 Add 12 to both sides of the equation 5y = 25 divide both sides by 5... y = 5 So we have x = -3 and y = 5

OpenStudy (anonymous):

@iambatman 4x+5y=13 3x+y=-4

OpenStudy (johnweldon1993):

Let me know if that doesn't make sense @dxgx3

OpenStudy (anonymous):

@johnweldon1993 THANK YOU SO MUCH !!!!